# Use of Nash equilibrium in finding game theoretic robust security bound on quantum bit error rate

### v1

@@ II Preliminaries of Game Theory
In this context, it is evident that the payoffs for Alice and Bob, as denoted by Eq. (2), are identical. For our subsequent analysis, we will treat these individual payoffs independently. This approach is taken due to the distinct probabilities, represented by $q$ for Alice and $p$ for Bob, pertaining to the selection of encoded bit values and the utilization of the $Z$ basis to prepare initial states.
@@ III Investigating Threshold Bound of QBER using Nash Equilibrium
A Nash equilibrium is a set of actions in a game where no player can improve their expected outcome (payoff) by altering their choice unilaterally, assuming the other players’ decisions remain unchanged. In a generalized concept of Nash equilibrium that represents a stochastic steady state of a strategic game, each player has the option to select a probability distribution over their available actions instead of being limited to a single, fixed choice. This probability distribution is referred to as a mixed strategy. In a well-defined game, it is generally assumed that all participants act logically and rationally. As a result, the primary goal for all players is to optimize (maximize in our case ) their expected payoffs. For simplicity, we evaluate mixed strategy Nash equilibrium [^3] A mixed strategy Nash equilibrium in a normal-form game is a set of mixed strategies for each player, where no player has an incentive to unilaterally deviate given the strategies chosen by others, ensuring mutual optimality. by taking three sets of subgames: $E_{1}$-$E_{2}$, $E_{1}$-$E_{3}$, and $E_{2}$-$E_{3}$ subgame [^4] All payoff elements for all parties under the attack scenarios $E_{1}$, $E_{2}$, $E_{3}$, and $E_{4}$ are computed in Appendices A, B, C, and D, respectively..
Firstly, we elucidate the best response function and subsequently apply it to various subgame scenarios. For each subgame, we will establish the best response function for each party. The best response for an individual is defined as the probability of choosing their rational choice in a mixed strategy scenario, aiming to achieve the best utility when other parties make their decisions independently and arbitrarily. Let’s consider the best response function for Alice in the subgame $E_{i}$ $E_{j}$. Suppose Alice has the probability $q$ of choosing her classical bit information when Bob and Eve independently choose their decisions with probabilities $p$ and [^5] Here, $r$ and $1-r$ denote the probabilities assigned to the selection of attacks $E_{i}$ and $E_{j}$ by an eavesdropper, Eve, respectively. $r$. For the pure strategy best response function of Alice, denoted as $q=1$ and $q=0$, it can be expressed as:
@@
$E_{1}$-$E_{2}$ subgame Previously defined probabilities $p$ and $q$ can be used. Let’s assume that Eve chooses the $E_{1}$($E_{2}$) attack with probability $r$ ($1-r$). We denote the best response function of Alice as $B_{A}\left(p,r\right)$, which represents the set of probabilities that Alice assigns to choosing the Z basis in her best responses to $p$ and $r$. Similarly, the best response functions for Bob and Eve are $B_{B}\left(q,r\right)$ and $B_{E}\left(p,q\right)$, respectively. We have,
@@
The best response functions for Alice, Bob, and Eve in the subgame scenarios $E_{1}$-$E_{3}$ and $E_{2}$-$E_{3}$ are as follows,
@@
The best response functions of various subgames are depicted in Figure 1. The mixed strategy Nash equilibrium in the game can be thought of as the points where the best response functions in the figure intersect. It is noteworthy that multiple points of intersection exist in different subgame scenarios, which signify the Nash equilibrium points for each player in these games. We provide a summary of this data, including the payoffs of each player, in Table 1 in Appendix E. In the provided summary, we analyze our mixed strategy Nash equilibrium points for different subgames and investigate the secure QBER bound. In the context of Eq. (2), it becomes evident that the payoff functions for the three parties depend on the variables $p,q$ ($q$) for Eve’s attack strategies, $E_{1}$, $E_{2}$ ( $E_{3}$, $E_{4}$). Furthermore, our game can be represented as a normal form game (strategic form game) because players have no information at their decision points about other players’ choices when making their moves. In this setup, all parties independently select their strategies. Specifically, Bob selects the $Z$ basis with a probability of $p$, Eve chooses the attack strategy with a probability of $r$, and Alice, in message mode, selects the encoded bit 00 with a probability of $q$. In our analysis, both Alice and Bob share the same payoff functions (2)). In the $E_{1}$-$E_{2}$ subgame, the expected payoffs for Alice/Bob and Eve are given by $rP_{A/B}^{E_{1}}(p,q)+(1-r)P_{A/B}^{E_{2}}(p,q)$ and $rP_{E}^{E_{1}}(p,q)+(1-r)P_{E}^{E_{2}}(p,q)$, respectively. By comparing the payoff differences between Eve and Alice (see in Table 1), we can identify among the Nash equilibrium points where Eve or Alice/Bob benefit the most. In the $E_{1}$-$E_{2}$ subgame, Eve benefits individually at the Nash equilibrium point $(0.45,0.195,0.005)$, while Alice/Bob benefits individually at the Nash equilibrium point $(0.72,0.208,0.225)$. Similarly, in the $E_{1}$-$E_{3}$ subgame, Eve and Alice/Bob benefit individually at Nash equilibrium points $\left(0.41,0.39,0.412\right)$ and $\left(0.84,0.047,0.525\right)$, respectively. In the $E_{2}$-$E_{3}$ subgame, Eve and Alice/Bob benefit individually at Nash equilibrium points $\left(0.385,0.215,0.262\right)$ and $\left(0.80,0.115,0.885\right)$, respectively. In conclusion, there is no Pareto optimal Nash equilibrium point in our subgames where payoffs would be favorable to the entire group of players.
@@
We calculate the QBER for all Nash equilibrium points in different subgames. It is pertinent to highlight that the expected QBER value for the $E_{i}$-$E_{j}$ subgame is defined as $\epsilon_{E_{i}-E_{2}}=r\,{\rm QBER}_{E_{i}}+\left(1-r\right){\rm QBER}_{E_{j}}$, where $r$ is a probability of Eve’s choice to perform $E_{i}$ attack. In any given subgame, Eve’s optimal scenario is characterized by either the maximum payoff difference (the discrepancy between Eve’s and Alice’s payoffs) or the minimum QBER. Our objective is to determine the QBER threshold value. Consequently, we seek the minimum QBER values across all subgame scenarios outlined in Table 1. These values correspond to the Nash equilibrium points, representing situations where Eve is strategically positioned to conceal her presence most effectively. The minimum QBER values at Nash equilibrium points are $0.610303$, $0.152451$, and $0.143882$ for the $E_{1}$-$E_{2}$, $E_{1}$-$E_{3}$, and $E_{2}$-$E_{3}$ subgames, respectively. The reduction in the minimum QBER value [^6] Considering lower QBER value empower Eve the most, that bound gives the most secure condition on a quantum protocol. (in message mode) suggests that more potent attack strategies by Eve are being applied in the subgame when the minimum or same $P_{d}$ (detection probability of Eve’s presence in control mode) is achieved for the attacks within the subgame. These lower QBER values are designed to identify more sophisticated quantum attacks by Eve within the game-like scenario. Following an analysis of the threshold values of QBER, we conclude that $E_{3}$ represents the most powerful quantum attack, followed by $E_{2}$, and then $E_{1}$. This implies that whenever Eve employs a powerful attack, she attains a higher payoff, posing a threat to both Alice and Bob. To address this threat in our DL04 protocol game scenario, we set a lower minimum value for the QBER corresponding to this specific potent attack by Eve. The threshold is determined by evaluating all Nash equilibrium points within the subgames that encompass the mentioned attack.
@@
As previously noted, a player with limited access to quantum resources appears to be a classical player. In our analysis, the $E_{4}$ attack is considered a classical attack by Eve, implying she is a classical eavesdropper when performing the $E_{4}$ attack. We also compare the classical attack $E_{4}$ with the simplest quantum attack, $E_{1}$. Furthermore, we analyze the $E_{4}$ attack (IR attack) in comparison with the $E_{1}$ attack as a $E_{1}$-$E_{4}$ subgame. In this subgame, the minimum value of QBER is found to be $0.323478$, which is lower than $E_{1}$-$E_{2}$ subgame. Despite this fact, $E_{4}$ is considered a less powerful attack than $E_{1}$ because the $P_{d}$ value is higher for $E_{4}$ (37.5%) compared to that for $E_{1}$ (18.75%) in control mode.
@@
To determine the secure upper bound of the QBER, denoted as $\epsilon$, considering all attack scenarios in the entire game, we focus on the subgame involving potent attacks which is the $E_{2}$-$E_{3}$ subgame. In this context, the upper bound of QBER is found to be $\epsilon=0.143882$ because this is the minimum QBER value for the subgame scenario where the most potent attack $E_{3}$ is present. The lower bound of QBER is 00 because if Eve applies only the $E_{3}$ attack, no error will be detected in message mode (see Appendix C). Furthermore, the lower and upper bounds of the detection probability of Eve’s presence ($P_{d}$) in control mode are established at $0.1875$ and $0.375$, respectively. Consequently, we can deduce that the DL04 protocol demonstrates $\epsilon$-secure, under the set of collective and individual (IR) attacks, denoted as $E_{1}$, $E_{2}$, $E_{3}$, and $E_{4}$.
@@
Figure 1: (Color online) The plot displays Nash equilibrium points. The low-density layer, medium-density layer, and high-density layer correspond to the best response functions of Alice, Bob, and Eve, respectively: (a) $E_{1}$-$E_{2}$ subgame, (b) $E_{1}$-$E_{3}$ subgame, (c) $E_{2}$-$E_{3}$ subgame, (d) $E_{1}$-$E_{4}$ subgame.
@@ IV Discussion
In this paper, we propose a new security definition for quantum communication protocols in the context of collective attacks and IR attack using Nash equilibrium. Alternatively, this can be referred to as a game-theoretic security boundary against a set of collective attacks employing the concept of Nash equilibrium. Nash equilibrium points represent stable points or offer optimistic and rational probabilities for stakeholders to make decisions in a mixed strategic game setting. By utilizing these probabilities derived from Nash equilibrium points, a stable game situation can be established for all parties, allowing for the evaluation of various cryptography parameters, including the determination of the threshold value of QBER. In a real-world scenario where decisions are made independently by all parties, achieving a stable point is less likely and may lead to a more unstable situation. Therefore, the identification of the lowest QBER value from the stable game scenario establishes a secure threshold boundary for the realistic implementation of the quantum protocol. Our analysis illustrates that the security (denoted as $\epsilon$-secure) of any quantum communication protocol depends on the choices of attack strategies employed by the eavesdropper, Eve. A smaller value of $\epsilon$ indicates a more potent attack strategy by Eve. Furthermore, our findings suggest that a quantum Eve possesses greater power than a classical Eve [^7] Ensuring that Eve’s selection of unitary operation and the execution of the measurement operation are appropriately synchronized within the quantum operation.. In our investigation, we assume that Eve has unlimited quantum resources, leading us to neglect the last three terms of Eve’s payoff in Eq. (1). To simplify our analysis, we also assume equal contributions of all payoff elements for each party, with weighted values set at $\frac{1}{4}$. We divide our game into four subgames and determine the Nash equilibrium points for each subgame. Subsequently, we identify the minimum value of QBER within each subgame, which corresponds to the secure bound of QBER for that specific subgame. Finally, we determine the threshold bound of QBER ($\epsilon$) for the entire game by evaluating the minimum QBER value within the subgame that encompasses the most potent attacks. Moreover, it has been observed that the DL04 protocol is vulnerable to Pavičić attacks in message mode due to the QBER being 00. In this scenario, security is ensured through the control mode executed jointly by both legitimate parties. However, the security concern for this attack will be mitigated through the control mode.
@@
It is noteworthy that quantum communication protocols serve various purposes, with security sometimes prioritized over efficiency, and vice versa. To address such trade-off scenarios, one may generalize our analysis by introducing $\omega_{i},\,\omega_{j},\,\omega_{k}\neq 0$, and $\omega_{m}\neq\omega_{n}$ if $m\neq n$. This implies distinct weighted values with each payoff, with the sum normalized to $1$ (ensuring weights assigned to players are normalized). Assuming $\omega_{i},\,\omega_{j},\,\omega_{k}\neq 0$, Eq. (1) yields $n_{1}=1$, $n_{2}=1$, $n_{3}=1$ for $E_{1}$, $n_{1}=4$, $n_{2}=2$, $n_{3}=1$ for $E_{2}$, $n_{1}=2$, $n_{2}=5$, $n_{3}=0$ for $E_{3}$, and $n_{1}=2$, $n_{2}=n_{3}=0$ for $E_{4}$ attack strategies. It is worth noting that the cost of applying multi-qubit gates increases with the rise in qubit numbers of that multi-qubit gate used. Importantly, our approach can be further generalized to encompass an entire game-like scenario with all these attacks together rather than a set of subgames. Under this assumption, different probabilities would exist for applying various attacks by eavesdroppers, leading to the identification of Pareto optimal Nash equilibrium points under specific conditions. Consequently, the expression for the QBER bound ($\epsilon$) becomes dependent on these generalized parameters.

### v2

@@ II Preliminaries of Game Theory
In this context, it is evident that the payoffs for Alice and Bob, as denoted by Eq. (2), are identical. For our subsequent analysis, we will treat these individual payoffs independently. This approach is taken due to the distinct probabilities, represented by $q$ for Alice and $p$ for Bob, pertaining to the selection of encoded bit values and the utilization of the $Z$ basis to prepare initial states. This is illustrated more clearly in the tabular format of the game $\mathcal{M}$ in Table 1.
Bob
Alice $Z\left(p\right)$ $X\left(1-p\right)$
$\left(rP_{A}^{E_{i}}\left(p,q\right)+\left(1-r\right)P_{A}^{E_{j}}\left(p,q\right),\right.$ $\left(rP_{A}^{E_{i}}\left(1-p,q\right)+\left(1-r\right)P_{A}^{E_{j}}\left(1-p,q\right),rP_{B}^{E_{i}}\left(1-p,q\right)\right.$
$0\left(q\right)$ $\left.rP_{B}^{E_{i}}\left(p,q\right)+\left(1-r\right)P_{B}^{E_{j}}\left(p,q\right),P_{E}^{E_{i/j}}\left(p,q\right)\right)$ $\left.+\left(1-r\right)P_{B}^{E_{j}}\left(1-p,q\right),P_{E}^{E_{i/j}}\left(1-p,q\right)\right)$
$1\left(1-q\right)$ $\left(rP_{A}^{E_{i}}\left(p,1-q\right)+\left(1-r\right)P_{A}^{E_{j}}\left(p,1-q\right),\right.$ $\left(rP_{A}^{E_{i}}\left(1-p,1-q\right)+\left(1-r\right)P_{A}^{E_{j}}\left(1-p,1-q\right),\right.$
$rP_{B}^{E_{i}}\left(p,1-q\right)+\left(1-r\right)P_{B}^{E_{j}}\left(p,1-q\right),$ $rP_{B}^{E_{i}}\left(1-p,1-q\right)+\left(1-r\right)P_{B}^{E_{j}}\left(1-p,1-q\right),$
$\left.P_{E}^{E_{i/j}}\left(p,1-q\right)\right)$ $\left.P_{E}^{E_{i/j}}\left(1-p,1-q\right)\right)$
Table 1: Payoff matrix for the $\mathcal{M}$ game. The first, second and third entry in the parenthesis denotes the payoff of Alice, Bob and Eve. Here, $r$ and $1-r$ denote the probabilities assigned to the selection of attacks $E_{i}$ and $E_{j}$ by an eavesdropper, Eve, respectively.
@@ III Investigating Threshold Bound of QBER using Nash Equilibrium
A Nash equilibrium is a set of actions in a game where no player can improve their expected outcome (payoff) by altering their choice unilaterally, assuming the other players’ decisions remain unchanged. In a generalized concept of Nash equilibrium that represents a stochastic steady state of a strategic game, each player has the option to select a probability distribution over their available actions instead of being limited to a single, fixed choice. This probability distribution is referred to as a mixed strategy. In a well-defined game, it is generally assumed that all participants act logically and rationally. As a result, the primary goal for all players is to optimize (maximize in our case ) their expected payoffs. For simplicity, we evaluate mixed strategy Nash equilibrium [^5] A mixed strategy Nash equilibrium in a normal-form game is a set of mixed strategies for each player, where no player has an incentive to unilaterally deviate given the strategies chosen by others, ensuring mutual optimality. by taking three sets of game scenarios $E_{1}$-$E_{2}$, $E_{1}$-$E_{3}$ and $E_{2}$-$E_{3}$ scenario where each player is looking to play a mixed quantum strategy that makes her opponent indifferent between her pure quantum strategies [^6] All payoff elements for all parties under the attack scenarios $E_{1}$, $E_{2}$, $E_{3}$ and $E_{4}$ are computed in Appendices A, B, C and D, respectively.. Based on this assumption, we can compare the results derived from Nash equilibrium points to obtain the secure bound of QBER (see Appendix E).
Firstly, we elucidate the best response function and subsequently apply it to various game scenarios. For each game scenario, we will establish the best response function for each party. The best response for an individual is defined as the probability of choosing their rational choice in a mixed strategy scenario, aiming to achieve the best utility when other parties make their decisions independently and arbitrarily. Let’s consider the best response function for Alice in the game $E_{i}$ $E_{j}$. Suppose Alice has the probability $q$ of choosing her classical bit information when Bob and Eve independently choose their decisions with probabilities $p$ and [^7] Here, $r$ and $1-r$ denote the probabilities assigned to the selection of attacks $E_{i}$ and $E_{j}$ by an eavesdropper, Eve, respectively. $r$. For the pure strategy best response function of Alice, denoted as $q=1$ and $q=0$, it can be expressed as [76]:
@@
$E_{1}$-$E_{2}$ game scenario Previously defined probabilities $p$ and $q$ can be used. Let’s assume that Eve chooses the $E_{1}$($E_{2}$) attack with probability $r$ ($1-r$). We denote the best response function of Alice as $B_{A}\left(p,r\right)$, which represents the set of probabilities that Alice assigns to choosing the Z basis in her best responses to $p$ and $r$ [76]. Similarly, the best response functions for Bob and Eve are $B_{B}\left(q,r\right)$ and $B_{E}\left(p,q\right)$, respectively. We have,
@@
The best response functions for Alice, Bob and Eve in the game scenarios $E_{1}$-$E_{3}$ and $E_{2}$-$E_{3}$ are as follows,
@@
In the context of a quantum game, Figure 2 shows the best response functions for different game scenarios. These functions represent each player’s optimal strategy in response to the strategies of the other players. The mixed strategy Nash equilibrium is found at the points where these best response functions intersect. This means that at these intersections, each player’s strategy is optimal given the strategies of the others. It is important to note that there can be multiple intersection points, each corresponding to different game scenarios. These points of intersection indicate the Nash equilibrium for each player in these game scenarios. In subfigures (a), (b), (c) and (d) of Figure 2, we identify the Nash equilibrium points for the $E_{1}$-$E_{2}$, $E_{1}$-$E_{3}$, $E_{2}$-$E_{3}$ and $E_{1}$-$E_{4}$ game scenarios, respectively. Each subfigure depicts three different curves, each representing the best response function of a player within the mixed quantum strategy game. These curves show the optimal strategies for each player in response to the others. The intersection points of these three curves indicate the most stable points in each game scenario. From these intersections, we can determine the most stable probability distributions in the mixed quantum strategy game, which are optimal for all players.
@@
We provide a summary of this data, including the payoffs of each player, in Table 2 in Appendix E. In the provided summary, we analyze our mixed strategy Nash equilibrium points for different game scenarios and investigate the secure QBER bound. In the context of Eq. (2), it becomes evident that the payoff functions for the three parties depend on the variables $p,q$ ($q$) for Eve’s attack strategies, $E_{1}$, $E_{2}$ ( $E_{3}$, $E_{4}$). Furthermore, our game can be represented as a normal form game (strategic form game) because players have no information at their decision points about other players’ choices when making their moves. In this setup, all parties independently select their strategies. Specifically, Bob selects the $Z$ basis with a probability of $p$, Eve chooses the attack strategy with a probability of $r$, and Alice, in message mode, selects the encoded bit 00 with a probability of $q$. In our analysis, both Alice and Bob share the same payoff functions (2)). In the $E_{1}$-$E_{2}$ game scenario, the expected payoffs for Alice/Bob and Eve are given by $rP_{A/B}^{E_{1}}(p,q)+(1-r)P_{A/B}^{E_{2}}(p,q)$ and $rP_{E}^{E_{1}}(p,q)+(1-r)P_{E}^{E_{2}}(p,q)$, respectively. By comparing the payoff differences between Eve and Alice (see in Table 2), we can identify among the Nash equilibrium points where Eve or Alice/Bob benefit the most. In the $E_{1}$-$E_{2}$ game scenario, Eve benefits individually at the Nash equilibrium point $(0.45,0.195,0.005)$, while Alice/Bob benefits individually at the Nash equilibrium point $(0.72,0.208,0.225)$. Similarly, in the $E_{1}$-$E_{3}$ game scenario, Eve and Alice/Bob benefit individually at Nash equilibrium points $\left(0.41,0.39,0.412\right)$ and $\left(0.84,0.047,0.525\right)$, respectively. In the $E_{2}$-$E_{3}$ game scenario, Eve and Alice/Bob benefit individually at Nash equilibrium points $\left(0.385,0.215,0.262\right)$ and $\left(0.80,0.115,0.885\right)$, respectively. In conclusion, there is no Pareto optimal Nash equilibrium point in our game scenarios where payoffs would be favorable to the entire group of players.
@@
We calculate the QBER for all Nash equilibrium points in different game scenarios. It is pertinent to highlight that the expected QBER value for the $E_{i}$-$E_{j}$ game scenario is defined as $\epsilon_{E_{i}-E_{2}}=r\,{\rm QBER}_{E_{i}}+\left(1-r\right){\rm QBER}_{E_{j}}$, where $r$ is a probability of Eve’s choice to perform $E_{i}$ attack. In any given game scenario, Eve’s optimal scenario is characterized by either the maximum payoff difference (the discrepancy between Eve’s and Alice’s payoffs) or the minimum QBER. Our objective is to determine the QBER threshold value. Consequently, we seek the minimum QBER values across all game scenarios outlined in Table 2. These values correspond to the Nash equilibrium points, representing situations where Eve is strategically positioned to conceal her presence most effectively. The minimum QBER values at Nash equilibrium points are $0.610303$, $0.152451$ and $0.143882$ for the $E_{1}$-$E_{2}$, $E_{1}$-$E_{3}$ and $E_{2}$-$E_{3}$ game scenarios, respectively. The reduction in the minimum QBER value [^8] Considering lower QBER value empower Eve the most, that bound gives the most secure condition on a quantum protocol. (in message mode) suggests that more potent attack strategies by Eve are being applied in this game scenario when the minimum or same $P_{d}$ (detection probability of Eve’s presence in control mode) is achieved for the attacks within this game scenario. These lower QBER values are designed to identify more sophisticated quantum attacks by Eve within the game-like scenario. Following an analysis of the threshold values of QBER, we conclude that $E_{3}$ represents the most powerful quantum attack, followed by $E_{2}$, and then $E_{1}$. This implies that whenever Eve employs a powerful attack, she attains a higher payoff, posing a threat to both Alice and Bob. To address this threat in our DL04 protocol game scenario, we set a lower minimum value for the QBER corresponding to this specific potent attack by Eve. The threshold is determined by evaluating all Nash equilibrium points within these game scenarios that encompass the mentioned attack.
@@
As previously noted, a player with limited access to quantum resources appears to be a classical player. In our analysis, the $E_{4}$ attack is considered a classical attack by Eve, implying she is a classical eavesdropper when performing the $E_{4}$ attack. We also compare the classical attack $E_{4}$ with the simplest quantum attack, $E_{1}$. Furthermore, we analyze the $E_{4}$ attack (IR attack) in comparison with the $E_{1}$ attack as a $E_{1}$-$E_{4}$ game scenario. In this game scenario, the minimum value of QBER is found to be $0.323478$, which is lower than $E_{1}$-$E_{2}$ game scenario. Despite this fact, $E_{4}$ is considered a less powerful attack than $E_{1}$ because the $P_{d}$ value is higher for $E_{4}$ (37.5%) compared to that for $E_{1}$ (18.75%) in control mode.
@@
To determine the secure upper bound of the QBER, denoted as $\epsilon$, considering all attack scenarios in the entire game, we focus on this game scenario involving potent attacks which is the $E_{2}$-$E_{3}$ game scenario. In this context, the upper bound of QBER is found to be $\epsilon=0.143882$ because this is the minimum QBER value for the game scenario where the most potent attack $E_{3}$ is present. The lower bound of QBER is 00 because if Eve applies only the $E_{3}$ attack, no error will be detected in message mode (see Appendix C). Furthermore, the lower and upper bounds of the detection probability of Eve’s presence ($P_{d}$) in control mode are established at $0.1875$ and $0.375$, respectively. Consequently, we can deduce that the DL04 protocol demonstrates $\epsilon$-secure, under the set of collective and individual (IR) attacks, denoted as $E_{1}$, $E_{2}$, $E_{3}$ and $E_{4}$.
@@
Figure 2: (Color online) The plot shows Nash equilibrium points where the best response functions of the three players intersect in our mixed strategy game scenario. The low-density layer, medium-density layer, and high-density layer correspond to the best response functions of Alice, Bob, and Eve, respectively: (a) $E_{1}$-$E_{2}$ game scenario, (b) $E_{1}$-$E_{3}$ game scenario, (c) $E_{2}$-$E_{3}$ game scenario, (d) $E_{1}$-$E_{4}$ game scenario.
@@ IV Discussion
In this paper, we propose a new security definition for quantum communication protocols in the context of collective attacks and IR attack using Nash equilibrium. Alternatively, this can be viewed as a game-theoretic security bound against collective attacks using Nash equilibrium. Nash equilibrium points are stable points that provide rational probabilities for decision-making in a mixed quantum strategy game. These probabilities create a stable game situation for all parties, enabling the evaluation of cryptographic parameters such as the threshold value of QBER. In a real-world scenario where decisions are made independently by all parties, achieving a stable point is less likely and may lead to a more unstable situation. Therefore, the identification of the lowest QBER value from the stable game scenario establishes a secure threshold boundary for the realistic implementation of the quantum protocol. Our analysis illustrates that the security (denoted as $\epsilon$-secure) of any quantum communication protocol depends on the choices of attack strategies employed by the eavesdropper, Eve. A smaller value of $\epsilon$ indicates a more potent attack strategy by Eve. Furthermore, our findings suggest that a quantum Eve possesses greater power than a classical Eve [^9] Ensuring that Eve’s selection of unitary operation and the execution of the measurement operation are appropriately synchronized within the quantum operation.. In our investigation, we assume that Eve has unlimited quantum resources, leading us to neglect the last three terms of Eve’s payoff in Eq. (1). To simplify our analysis, we also assume equal contributions of all payoff elements for each party, with weighted values set at $\frac{1}{4}$. We divide our game into four different game scenarios and determine the Nash equilibrium points for each game scenario. Subsequently, we identify the minimum value of QBER within each game scenario, which corresponds to the secure bound of QBER for that specific game scenario. Finally, we determine the threshold bound of QBER ($\epsilon$) for the entire game by evaluating the minimum QBER value within the game scenario that encompasses the most potent attacks. Moreover, it has been observed that the DL04 protocol is vulnerable to Pavičić attacks in message mode due to the QBER being 00. In this scenario, security is ensured through the control mode executed jointly by both legitimate parties. However, the security concern for this attack will be mitigated through the control mode.
@@
It is noteworthy that quantum communication protocols serve various purposes, with security sometimes prioritized over efficiency, and vice versa. To address such trade-off scenarios, one may generalize our analysis by introducing $\omega_{i},\,\omega_{j},\,\omega_{k}\neq 0$, and $\omega_{m}\neq\omega_{n}$ if $m\neq n$. This implies distinct weighted values with each payoff, with the sum normalized to $1$ (ensuring weights assigned to players are normalized). Assuming $\omega_{i},\,\omega_{j},\,\omega_{k}\neq 0$, Eq. (1) yields $n_{1}=1$, $n_{2}=1$, $n_{3}=1$ for $E_{1}$, $n_{1}=4$, $n_{2}=2$, $n_{3}=1$ for $E_{2}$, $n_{1}=2$, $n_{2}=5$, $n_{3}=0$ for $E_{3}$ and $n_{1}=2$, $n_{2}=n_{3}=0$ for $E_{4}$ attack strategies. It is important to note that the cost of using multi-qubit gates goes up as the number of qubits in the gate increases. Importantly, our approach can be further generalized to encompass an entire game-like scenario with all these attacks together rather than a set of different game scenarios. Under this assumption, different probabilities would exist for applying various attacks by eavesdroppers, leading to the identification of Pareto optimal Nash equilibrium points under specific conditions. Consequently, the expression for the QBER bound ($\epsilon$) becomes dependent on these generalized parameters.